### v1

# ASGO: Adaptive Structured Gradient Optimization

Abstract
Training deep neural networks (DNNs) is a structured optimization problem, because the parameters are naturally represented by matrices and tensors rather than simple vectors. Under this structural representation, it has been widely observed that gradients are low-rank and Hessians are approximately block-wise diagonal. These structured properties are crucial for designing efficient optimization algorithms but may not be utilized by current popular optimizers like Adam. In this paper, we present a novel optimization algorithm ASGO that capitalizes on these properties by employing a preconditioner that is adaptively updated using structured gradients. By fine-grained theoretical analysis, ASGO is proven to achieve superior convergence rates compared to existing structured gradient methods. Based on the convergence theory, we further demonstrate that ASGO can benefit from the low-rank and block-wise diagonal properties. We also discuss practical modifications of ASGO and empirically verify the effectiveness of the algorithm on language model tasks.
1 Introduction
Numerical optimization algorithms, especially those can efficiently train large foundation models (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023, ; Ouyang et al., 2022), play an important role in the modern machine learning field. Among them, adaptive gradient methods like AdaGrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2014) are popular choices, gaining huge success in training state-of-the-art models in many tasks. These algorithms typically uses a diagonal matrix preconditioner in the update, like for AdaGrad
$w_{t+1}=w_{t}-\eta_{t}\Lambda_{t}^{-1}g_{t},\text{ where $w_{t}\in\mathbb{R}^{d}$, $g_{t}\in\mathbb{R}^{d}$, and $\Lambda_{t}\in\mathbb{R}^{d\times d}$ is a diagonal matrix}.$
This coordinate-wise step size design has been theoretically verified to be effective as it can exploit the sparsity of the gradient vectors $g_{t}$ (Duchi et al., 2011). Also, when the Hessian is well-approximated by a diagonal matrix whose diagonal entries have very different scales, adaptive gradient methods have been proven to be beneficial (Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2024). While these results seem to be convincing, common deep neural networks (DNNs) do not necessarily have sparse gradients or Hessians that are well-approximated by ill-conditioned diagonal matrices. Instead, if we take the matrix structure of gradients in neural networks into account, it has been widely observed that these structured gradients are usually low-rank (Zhao et al., 2021; Yang et al., 2023; Cosson et al., 2023), and the Hessians are well-approximated by block-wise diagonal matrices (Collobert, 2004; Zhang et al., 2024a, b). Since adaptive gradient methods simply treat the parameters as vectors and ignore the matrix structure of gradients, they are generally unable to exploit these structured properties. This gap makes us ask the following question: How can we properly consider the matrix structures of gradients and exploit the low-rank and block-wise diagonal properties?
One possible attempt is Shampoo (Gupta et al., 2018), which considers the matrix (or tensor) structure of gradients:
$W_{t+1}=W_{t}-\eta_{t}L_{t}^{-\frac{1}{4}}G_{t}R_{t}^{-\frac{1}{4}},\text{ where $W_{t}\in\mathbb{R}^{m\times n}$ and $L_{t}\in\mathbb{R}^{m\times m}$, $R_{t}\in\mathbb{R}^{n\times n}$ are full matrices.}$.
The main motivation for such design is that if we apply vectorization for the update, the Shampoo preconditioner is a single matrix that is the Kronecker product of $L_{t}$ and $R_{t}$, which is an approximation of the full-matrix preconditioner for AdaGrad (Duchi et al., 2011). However, the theoretical convergence of Shampoo is worse than that for AdaGrad or even SGD when the dimension is large. Also, Shampoo needs more memory and much heavier computation than adaptive gradient methods because it requires two preconditioners, making it less suitable for training large-scale DNNs.
In this paper, we provide an answer to the aforementioned question by proposing ASGO (Adaptive Structured Gradient Optimization), which significantly improves the convergence guarantees of Shampoo while requiring less memory and computation. In light of the analysis that demonstrates the benefits of adaptive gradient methods (Duchi et al., 2011; Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2020), we use appropriate assumptions to enable a more fine-grained convergence analysis, showing superior convergence results and how ASGO can benefit from the low-rank and block-wise diagonal properties of the problem. We also discuss the connection between ASGO and Muon (Jordan et al., 2024), a structured gradient method originated from the steepest descent algorithm with spectral norm, to conjecture a relation between ASGO and Muon analogous to the relation between AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). Furthermore, we develop an efficient design for query-key attention parameters in transformer models and examine the empirical performance on pretraining transformer model tasks.
Our main contributions are summarized as follows:
  • We propose ASGO, an algorithm utilizing the structured gradient properties. We theoretically analyze the convergence of ASGO and show that it converges faster than full-matrix AdaGrad and Shampoo.
  • We further demonstrate that ASGO can effectively exploit the low-rankness of gradients as well as the approximate block-wise diagonal property of Hessians that is typically observed in training DNNS, hence indicating that ASGO has a great potential for real-world applications.
  • We develop a practical implementation of ASGO with targeted modifications for transformer architectures that offers significant advantages: it eliminates the need for a separate optimizer for 1D parameters like Muon, while requiring less memory and computational complexity compared to Shampoo.
  • We empirically validate the effectiveness of ASGO and the modifications on Transformer pretraining tasks, demonstrating the algorithm’s great potential in deep learning tasks.
2 Related Work
Adaptive Gradient Methods.
Adaptive gradient methods that use diagonal preconditioners to speedup the convergence are extremely popular for solving many real-world optimization problems. To the best of our knowledge, the first method of this kind for machine learning, AdaGrad (Duchi et al., 2011; Streeter and McMahan, 2010), was developed based on rigorous theory that showed the benefits of using this kind of preconditioner. Adam (Kingma and Ba, 2014; Loshchilov and Hutter, 2017) modified AdaGrad and has become the default choice for training large foundation models. In theory, it has been proven that adaptive gradient methods can benefit from sparse gradients and approximately ill-conditioned Hessians (Duchi et al., 2011; Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2024). It is worth noting that the original AdaGrad paper (Duchi et al., 2011) also proposed a version of AdaGrad that uses a full-matrix preconditioner instead of the diagonal one, which is believed to perform even better. However, this full-matrix AdaGrad method suffers from large memory costs for storing the preconditioner, and there is no better convergence guarantee compared to diagonal AdaGrad and SGD under the same settings as listed in Section 4.
Optimization with Matrix Structure.
In the standard optimization literature (i.e., excluding areas such as conic optimization), it is common to consider variables as vectors. However, recently, optimization methods for machine learning that consider variables as matrices have been rapidly gaining attention. Adafactor (Shazeer and Stern, 2018), LAMB (You et al., 2019), and Adam-mini (Zhang et al., 2024b) consider the matrix or layer structure to help reduce the memory cost of Adam and enable more efficient training. KFAC (Martens and Grosse, 2015) and Shampoo (Gupta et al., 2018) are two pioneering works that approximate, respectively, the full-matrix preconditioner of AdaGrad and the empirical Fisher matrix, using Kronecker products of smaller matrices, making the memory cost more affordable. Unfortunately, the analysis in (Gupta et al., 2018) shows that the rate of convergence for Shampoo is no better than that for the full-matrix AdaGrad. We note that there is another method, TNT (Ren and Goldfarb, 2021) that is very closely related to Shampoo. TNT was developed as a natural gradient method, approximating the true Fisher matrix (FM) by the covariance of block-wise sampling-based gradients assuming that they are Tensor-Normally distributed. The main difference between TNT and Shampoo, besides that between the FM and EF matrices, is that TNT uses the inverse of their Kronecker factors, whereas Shampoo uses the -1/2 power of them.
Many recent works have focused on improving full-matrix AdaGrad and Shampoo. Feinberg et al. (2023) uses a sketching-based approach to approximate the full-matrix AdaGrad preconditioner with lower memory cost. Morwani et al. (2024) provides theoretical intuition and empirical evidence to claim that Shampoo should use the $-1/2$ power on the preconditioners to better approximate the EF matrix. Vyas et al. (2024) demonstrates that Shampoo is like doing Adafactor in the eigenspace of gradients and proposes a novel algorithm, SOAP, that performs Adam in this eigenspace. SOAP is observed to achieve better performance than Adam and Shampoo, but suffers from a high computation load per iteration. Galore (Zhao et al., 2024) shares a similar algorithmic design with SOAP with extra focus on lowering memory costs. Muon (Jordan et al., 2024) follows this line of work, using the steepest descent framework with the spectral norm, and is shown to be scalable and effective in training large foundation models (Liu et al., 2025a). More recently, Nguyen et al. (2025) proposes AdaDiag, which may be viewed as SOAP doing SVD without gradient accumulation. Liu et al. (2025b) proposes COSMOS, a combination of SOAP and Muon, that trades off between performance and computational efficiency.
Rank of Gradients and Weight Updates.
It has been widely observed that gradients are naturally low-rank in DNNs, even when a large batch size is employed (Gur-Ari et al., 2018; Zhao et al., 2021; Yang et al., 2023; Cosson et al., 2023). This property has been widely utilized for computation and memory efficiency in training (Wang et al., 2018; Cosson et al., 2023; Zhao et al., 2024). On the other hand, the rank of the total weight update $\Delta W=W_{T}-W_{0}$, actually depends a lot on the training method. If we use LoRA (Hu et al., 2022), $\Delta W$ is determined to be low-rank. However, in pretraining or even many complex fine-tuning tasks, LoRA is observed to perform much worse than methods that produce high-rank weight updates like full-parameter training, which is conjectured to be due to the weight update rank (Lialin et al., 2023; Jiang et al., 2024b; Huang et al., 2025).
Block-wise Diagonal Hessian.
It has been observed that the Hessian of a neural network tends to be block-wise diagonal with each block corresponding to a neuron both in experiments and theory for small MLPs (Collobert, 2004). Zhang et al. (2024a, b) recently numerically verified this property in small transformers and Zhang et al. (2024a) further empirically showed that transformers may exhibit heterogeneity between blocks, while CNNs may not.
Concurrent Work.
When we were finishing writing this paper, we noticed that concurrent work (Xie et al., 2025), which also investigates the same algorithm as ASGO, referred to as One-Sided Shampoo in their paper, just appeared on arXiv. Although the theoretical techniques and convergence results for ASGO are very similar in general, there are still some notable differences between the two works. First, the motivations are different. In Xie et al. (2025), the authors develop ASGO from the unified preconditioning method framework AdaptReg (Gupta et al., 2017) and mainly highlight its superior convergence results compared to Shampoo and full-matrix AdaGrad. In our paper, we focus more on theoretically discussing how ASGO can utilize the structured properties of optimization problems including low-rank gradients and block-wise diagonal Hessians, to highlight the potential of ASGO as a practical algorithm for training deep learning models. Second, our empirical results provide evidence that ASGO can perform well in practical tasks, whereas Xie et al. (2025) focuses on convex settings and examines One-Sided Shampoo only on linear regression. Moreover, our implementation includes specialized designs for Transformer architectures (particularly query/key) to improve performance on deep learning training tasks. As a direct byproduct of our main algorithm, we also developed a lightweight diagonal version named DASGO to trade off memory consumption and performance. Third, Xie et al. (2025) focuses primarily on developing a general proof framework applicable to multiple optimizers including One-Sided Shampoo, whereas our work specifically examines the theoretical and practical benefits of ASGO. To conclude, both works contribute to a better understanding of ASGO/One-Sided Shampoo.
3 Preliminaries and the Algorithm
3.1 Notation
In the following analysis, for an arbitrary matrix $W\in\mathbb{R}^{m\times n}$, we denote
  • $\left\|W\right\|_{\mathrm{op}}$ as the spectral norm of a matrix $W$, i.e., the largest singular value of it;
  • $\left\|W\right\|_{*}$ as the trace norm of a matrix $W$, i.e., the summation of its singular values, which is well-known as the dual norm of the spectral norm;
  • $\left\|W\right\|_{\mathrm{F}}$ as the Frobenius norm of $W$, which also equals ${\rm tr}\left(W^{\top}W\right)$, where ${\rm tr}\left(\cdot\right)$ is the trace;
  • $\left\|W\right\|_{L}\triangleq{\rm tr}\left(W^{\top}LW\right)$, where $L\in\mathbb{R}^{m\times m}$ is a real symmetric positive definite matrix.
We use ,,$\prec$,$\preceq$,$\succ$,$\succeq$ , ⪯ , ≻ , ⪰ to denote the Lowner order (Löwner, 1934) of square matrices, i.e., for matrices $A,B\in\mathbb{R}^{m\times m}$, $A\preceq B$ implies that $B-A$ is positive semidefinite and $A\prec B$ implies that $B-A$ is positive definite. $\succ$ and $\succeq$ are defined accordingly. Throughout this paper, we typically use capital letters such as $W$ to present matrices and $[W]_{i,j}$ to present the $(i,j)$-th entry of $W$.
3.2 Problem Settings
We study the following stochastic optimization problem:
$\min_{W\in\mathbb{R}^{m\times n}}f(W)\triangleq\mathbb{E}_{\xi}\left[f(W,\xi)\right],$ (1)
where we can only access to the stochastic gradient oracle $\nabla f(W;\xi)$ at $W$. In our analysis of convergence, we will mainly consider the convex optimization case.
Assumption 1 (Convexity).
$f(\cdot)$ is convex and $W_{*}$ is one of its minimizers.
Although deep neural networks are in general nonconvex globally, convexity may apply locally in some regions. Thus, convex analysis can be helpful for understanding the behavior of algorithms.
Algorithm 1 ASGO (Adaptive Structured Gradient Optimization)
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, $\epsilon\in\mathbb{R}$, schedule $\{\eta_{t}\}$ and batch size $M\in\mathbb{N}$ ($\epsilon$ should be small, similar to the $\epsilon$ for Adam or AdaGrad.)
2:  Initialize $V_{-1}=0\in\mathbb{R}^{m\times m}$
3:  for $t=0$ to $T-1$ do
4:     Sample mini-batch $\mathcal{B}_{t}$ with $\left|\mathcal{B}_{t}\right|\equiv M$ uniformly
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$triangleright Compute stochastic (sub)gradient
6:     $V_{t}=V_{t-1}+G_{t}G_{t}^{\top}$
7:     $\Lambda_{t}=V_{t}^{\frac{1}{2}}+\epsilon I_{m}$triangleright Use the square root of the summation to update
8:     $W_{t+1}=W_{t}-\eta_{t}\Lambda_{t}^{-1}G_{t}$
9:  end for
3.3 ASGO (Algorithm 1)
We propose ASGO (Adaptive Structured Gradient Optimization) in Algorithm 1, which is an algorithm with a single-side preconditioner. Compared to full-matrix AdaGrad, ASGO preserves and utilizes the matrix structure of $W_{t}$ and $G_{t}$, avoiding the huge memory cost for the storage of its preconditioner. Compared to Shampoo, ASGO employs only single preconditioner from one side, leading to the following main update rule:
$W_{t+1}=W_{t}-\eta_{t}V_{t}^{-\frac{1}{2}}G_{t},\text{ where $W_{t}\in\mathbb{R}^{m\times n}$ and $V_{t}\in\mathbb{R}^{m\times m}$ is a full matrix,}$,
if we ignore $\epsilon$, which is used for numerical stability. Basically, employing only one preconditioner largely reduces the memory cost and also improves the computational efficiency of Shampoo, as we only need to store one preconditioner matrix and only need to compute one matrix square root and inverse. On the other hand, this change biases from the starting point of Shampoo, as the preconditioner is no longer an effective approximation of the full-matrix AdaGrad preconditioner or the empirical Fisher matrix (Gupta et al., 2018; Morwani et al., 2024). However, as we will see in the following sections, this design can actually better exploit the low-rankness of gradients and the block-wise diagonal nature of Hessians, in terms of achieving better convergence rates.
4 Nonsmooth Theory
Theorem 1 (Nonsmooth convergence).
Under Assumption 1, for Algorithm 1 with $\eta_{t}\equiv\eta=D_{\mathrm{op}}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})]-f(W_{*})\leq\frac{1}{T}\mathbb{E}\left[\left\|\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]\cdot D_{\mathrm{op}}+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T},$
where $D_{\mathrm{op}}\triangleq\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{op}}$ and $D_{\mathrm{F}}\triangleq\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{F}}$.
Corollary 2.
If we additionally assume an upper bound for each stochastic (sub)gradient such that $\mathbb{E}\left[G_{t}G_{t}^{\top}\right]\preceq Q^{2}$, where $Q\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix, Theorem 1 also implies
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})]-f(W_{*})\leq\mathcal{O}\left(\frac{\left\|Q\right\|_{*}D_{\mathrm{op}}}{\sqrt{T}}+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}\right).$
Let us make some comments regarding the bound. First, one can easily check that the convergence rate should be in general $\mathcal{O}(1/\sqrt{T})$, which is standard for convex nonsmooth problems. An issue here might be that we treat $D_{\mathrm{op}}$ as a constant here, but it may increase as the iteration number $T$ increases. This can be addressed, for example, by involving a projection onto a bounded convex set $\mathcal{W}$, which should be with respect to the norm $\left\|\cdot\right\|_{\Lambda_{t}}$ in each iteration, just like AdaGrad does (Duchi et al., 2011). However, since this projection step is rarely used in training DNNs, we follow Gupta et al. (2018) to omit it in the algorithm to make it more practical. More importantly, this theoretical bound depends on the trace norm of gradients and the spectral norm of weights, showing that the algorithm can make use of the low-rank property of gradients. We would like to share some more points and intuitions about ASGO and the convergence theorem through the following discussions.
  • Comparison with SGD. SGD has the convergence rate of (Zinkevich, 2003):
[…]
    where $D_{\mathrm{F}}$ and $\left\|Q\right\|_{\mathrm{F}}$ indicates the Frobenius norm upper bound for weights and gradients respectively. By comparing this bound with Corollary 2,
    • we have $\left\|Q\right\|_{\mathrm{F}}\leq\left\|Q\right\|_{*}\leq\sqrt{r_{G}}\left\|Q\right\|_{\mathrm{F}}$, where $r_{G}$ is the rank of $Q$. Thus when $G_{t}$ are low-rank, or have very imbalanced singular values, $\left\|Q\right\|_{*}$ can be close to $\left\|Q\right\|_{\mathrm{F}}$;
    • we have $D_{\mathrm{F}}/\sqrt{r_{D}}\leq D_{\mathrm{op}}\leq D_{\mathrm{F}}$, where $r_{D}$ is the max rank of $W_{t}-W_{*}$. Thus, when $W_{t}-W_{*}$ are relatively high-rank or have lots of singular values in a similar scale, $D_{\mathrm{op}}$ can be much smaller than $D_{\mathrm{F}}$.
    Therefore, we can see in general, ASGO should work well when $G_{t}$ are low-rank and $W_{t}-W_{*}$ are relatively high-rank.
  • Intuitions on ASGO in practical tasks. We argue that low-rank $G_{t}$ and relatively high-rank $W_{t}-W_{*}$ should be common in many practical tasks, revealing the potential of ASGO in applications. As we have discussed in Section 2, gradients are commonly low-rank in deep neural networks, as verified by Zhao et al. (2021); Yang et al. (2023); Cosson et al. (2023). Meanwhile, given the huge success of LoRA (Hu et al., 2022) in fine-tuning foundation models, which naturally asserts a low-rank total update in $W$, there may seem to be a conflict to assume that $W_{t}-W_{*}$ has a high rank. However, it has been observed that $W_{0}-W_{*}$ should be relatively high-rank to obtain a better result, at least in pretraining and some complex fine-tuning tasks for large foundation models (Lialin et al., 2023; Jiang et al., 2024b; Huang et al., 2025). Further exploration of the connection between the rank of weight updates and the performance of algorithms will be an interesting topic.
  • Comparison with Full-Matrix AdaGrad and Shampoo. Shampoo (Gupta et al., 2018) and the full-matrix AdaGrad (Duchi et al., 2011) achieve the following convergence rates under the same settings as Theorem 1.
    Full-Matrix AdaGrad: $\quad\mathcal{O}\left(D_{\mathrm{F}}\sum_{j=1}^{m}\sum_{i=1}^{n}\sqrt{\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right)$
    Shampoo: $\quad\mathcal{O}\left(\sqrt{r}D_{\mathrm{F}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)^{\frac{1}{4}}\right)\right).$
    We can check that Theorem 1 indicates a convergence speed that is at least $D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than full-matrix AdaGrad and $\sqrt{r_{G}}D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than Shampoo (proofs are available in Appendix C). This also provides theoretical evidence that single-side preconditioning may better exploit the low-rankness of gradients, showing a faster convergence speed compared to Shampoo-like preconditioning.
Remark 1.
The proof is available in Appendix C. We keep the matrix structure of $W_{t}$ and $G_{t}$ throughout our analysis, in contrast with the standard analysis of the convergence of AdaGrad and Shampoo, which is based on vectorizations of $W_{t}$ and $G_{t}$. This is important for proving that ASGO can exploit the structured properties.
5 Smooth Theory
It is also important to study the performance of the algorithm in smooth settings, as many real training tasks have been widely observed to be smooth, at least locally. Also, only in smooth settings can we properly describe the importance of batch size.
Assumption 2 (Smoothness).
$f$ is $1$-smooth with respect to $\left\|\cdot\right\|_{L}$, where $L\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix and for any $X\in\mathbb{R}^{m\times n}$,
$\left\|X\right\|_{L}^{2}\triangleq{\rm tr}\left(X^{\top}LX\right).$
If we denote $X=[x_{1},\dots x_{n}]$, where each $x_{i}\in\mathbb{R}^{m}$, and vectorize $X$, we obtain that
${\rm tr}\left(X^{\top}LX\right)=\sum_{i=1}^{n}x_{i}^{\top}Lx_{i}=\begin{bmatrix}x_{1}\\ \cdots\\ x_{n}\end{bmatrix}^{\top}\begin{bmatrix}L&&&\\ &L&&\\ &&\cdots&\\ &&&L\end{bmatrix}\begin{bmatrix}x_{1}\\ \cdots\\ x_{n}\end{bmatrix},$
which means that Assumption 2 is equivalent to the existence of a symmetric matrix $L\in\mathbb{R}^{m\times m},L\succ 0$ such that for any $w\in\mathbb{R}^{mn}$,
$-\begin{bmatrix}L&&&\\ &L&&\\ &&\cdots&\\ &&&L\end{bmatrix}\preceq\nabla^{2}f_{v}(w)\preceq\begin{bmatrix}L&&&\\ &L&&\\ &&\cdots&\\ &&&L\end{bmatrix}\in\mathbb{R}^{mn\times mn},$
where $f_{v}(w)=f(W)$ if $W=[w_{1},\dots,w_{n}]\in\mathbb{R}^{m\times n}$ and $w=[w_{1};w_{2};\dots;w_{n}]\in\mathbb{R}^{mn}$ is the vectorization of $W$. Therefore, this assumption is closely related to the blockwise diagonal structure of the Hessian as discovered in practice (Collobert, 2004; Zhang et al., 2024a). Also note that Assumption 2 implies the standard smoothness assumption with respect to the Frobenius norm by $\left\|\nabla f_{v}(w)\right\|_{\mathrm{op}}\leq\left\|L\right\|_{\mathrm{op}}$. This block-wise diagonal smoothness assumption is a different direction for extending the smoothness condition from the diagonal anisotropic smoothness assumption used for sign-based and adaptive gradient methods (Bernstein et al., 2018; Liu et al., 2024).
Assumption 3 (Variance).
Let $N\triangleq\nabla f(W;\xi)-\nabla f(W_{t})\in\mathbb{R}^{m\times n}$ be the stochastic gradient noise. We assume that $\mathbb{E}[N]=0$ and there exists a symmetric positive definite matrix $V$ such that
$\mathbb{E}\left[NN^{\top}\right]\preceq V^{2}.$
One can check that Assumption 3 implies the standard variance bound by $\mathbb{E}[\left\|N_{t}\right\|_{\mathrm{F}}^{2}]\leq\left\|V\right\|_{\mathrm{F}}^{2}$. The assumption shares some similarity with the coordinate-wise variance bound as in Bernstein et al. (2018); Crawshaw et al. (2022); Liu et al. (2024), in the sense that it allows a more fine-grained analysis. This matrix-form variance upper bound may better describe the real case since it takes the structure of the noise into account, which is relevant to the rank and other structured properties.
(a) Hessian of a MLP (Collobert, 2004)
(b) Hessian of a MLP at initialization
(c) Hessian of a MLP at 50% step
(d) Hessian of a MLP at 100% step
Figure 1: This figure is from Zhang et al. (2024a, b), describing the near-block-wise diagonal Hessian structure of neural nets. (a) is the Hessian of an MLP after 1 training step reported in (Collobert, 2004). (b,c,d): the Hessians of a 1-hidden-layer MLP with 8 neurons on CIFAR-100. The near-block-diagonal structure maintains throughout training, where each block corresponds to one neuron.
Theorem 3 (Smooth Convergence).
Under Assumptions 1, 2 and 3, for Algorithm 1 with $\eta_{t}\equiv\eta=D_{\mathrm{op}}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]\leq\frac{4D_{\mathrm{op}}^{2}\left\|L\right\|_{*}}{T}+\frac{2\sqrt{2}D_{\mathrm{op}}\left\|V\right\|_{*}}{\sqrt{MT}}+\frac{2\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T},$
where $M$ is the batch size, $D_{\mathrm{op}}\triangleq\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{op}}$, and $D_{\mathrm{F}}\triangleq\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{F}}$..
As also discussed in Section 4, Theorem 3 depends on $D_{\mathrm{op}}$ and $D_{\mathrm{F}}$, which can be bounded if we additionally employ a projection step when doing the update. Theorem 3 indicates a similar general convergence rate to Corollary 2 for a small batch size, where the $\mathcal{O}(1/\sqrt{MT})$ term should be dominant in the rate, and thus shares the same properties as we discussed in Section 4. This means that ASGO can benefit if $V$, i.e., the low-rank stochastic gradient noise, is generally low-rank. Furthermore, when the batch size $M$ is large such that the $\mathcal{O}(\left\|L\right\|_{*}/T)$ term is also significant in the bound, Theorem 3 implies more, which can be reflected in the following discussions.
  • Comparison with SGD. SGD has the convergence rate of (Garrigos and Gower, 2023):
[…]
    where $D_{\mathrm{F}}$ indicates the Frobenius norm upper bound for weights. Since the comparison between the $\mathcal{O}(1/\sqrt{MT})$ term is generally consistent with the discussion in Section 4, we further compare the $\mathcal{O}(1/T)$ term here to provide some more intuitions:
    • we have $\left\|L\right\|_{\mathrm{op}}\leq\left\|L\right\|_{*}\leq r_{L}\left\|L\right\|_{\mathrm{op}}$, where $r_{L}$ is the rank of $L$. Thus when $L$ is low-rank, or have very imbalanced singular values, $\left\|L\right\|_{*}$ can be close to $\left\|L\right\|_{\mathrm{op}}$;
    • we have $D_{\mathrm{F}}/\sqrt{r_{D}}\leq D_{\mathrm{op}}\leq D_{\mathrm{F}}$, where $r_{D}$ is the max rank of $W_{t}-W_{*}$. Thus when $W_{t}-W_{*}$ are relatively high-rank or have lots of singular values in a similar scale, $D_{\mathrm{op}}$ can be much smaller than $D_{\mathrm{F}}$.
    Therefore, we can see in general, ASGO should work well when the Hessian can be well approximated by a block-wise diagonal matrix with a low-rank $L$ for each block and $W_{t}-W_{*}$ are relatively high-rank. Note that Hessians are discovered to be low-rank in DNNs, especially after some steps of training (Sagun et al., 2016, 2017; Wu et al., 2020).
  • Block-wise diagonal structure of Neural Networks. It has been widely observed that the Hessian of MLPs, as well as deep neural networks, are approximately block-wise diagonal (Collobert, 2004; Zhang et al., 2024a, b), as shown in Figure 1. This block-wise diagonal structure naturally arises from the structure of MLPs, i.e., the use of neurons. Inside each neuron, the parameters are more closely related and thus have a denser Hessian matrix, while outside the neuron, the Hessian can be very sparse. Based on Theorem 3, our algorithm should perform well in this block Hessian setting, showing great potential in real applications.
  • Intuitions on single-side preconditioner. Since we only use a single-side preconditioner for ASGO, it is also important to determine on which side we shall apply the preconditioner. As demonstrated in Figure 1, a block in the Hessian commonly corresponds to a neuron in the next layer, which is presented as a row of the weight matrix in MLPs. Therefore, we may desire the single-side preconditioner to be on the right side of $G_{t}$ to correctly fit this block-wise diagonal property. Based on this block-wise diagonal structure, if Hessian is approximately block-wise diagonal, Shampoo-like double-side preconditioners may not only suffer from higher memory and computation costs, but also a less effective preconditioning since they are trying to approximate the structured curvature information matrix using a dense matrix produced by Kronecker product (Gupta et al., 2018; Morwani et al., 2024).
Remark 2.
The proof of Theorem 3 is available in Appendix D. The analysis follows a similar mainline as the smooth analysis for AdaGrad (Levy et al., 2018; Liu et al., 2024), but is more complex because of the involvement of matrix operations in ASGO.
6 More Discussions on the Algorithm
Connection with Muon.
Muon (Jordan et al., 2024) can be interpreted as a standard steepest descent algorithm utilizing the spectral norm (Bernstein and Newhouse, 2024) with momentum if we ignore the incorporated momentum, i.e.
$W_{t+1}=$ $\underset{W\in\mathbb{R}^{m\times n}}{\text{argmin}}\left\{\left\langle G_{t},W-W_{t}\right\rangle+\frac{1}{2\eta_{t}}\left\|W-W_{t}\right\|_{\mathrm{op}}^{2}\right\},$ (2)
which preserves the matrix structure of the problem and should naturally exploit the structured properties because of the involvement of spectral norm in the steepest descent framework. This is consistent with what we discussed for ASGO. Moreover, if we ignore the gradient accumulation in the preconditioner and momentum, we can see that ASGO is equivalent to Muon. [^1] A proof can be found in Appendix E. Interestingly, this equivalence between ASGO and Muon in both theoretical intuition and algorithm design is analogous to that between diagonal AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). In this sense, we may interpret Muon and ASGO as SignSGD and AdaGrad for structured gradients, respectively. Note that Shampoo also admits such a relation with Muon in algorithmic design. However, as discussed in Section 4, Shampoo achieves a much worse convergence rate and thus fails to benefit from the structured properties as ASGO does. To some extent, this means that ASGO may be a more appropriate approach than Shampoo as such an analog to diagonal AdaGrad.
From this intuition, we may also conjecture the nonconvex convergence rate of ASGO. We prove the following non-convex convergence for Muon:
$\min_{0\leq t\leq T-1}\left\|\nabla f(W_{t})\right\|_{*}^{2}\leq\mathcal{O}\left(\frac{\left\|L\right\|_{*}(f(W_{0})-f^{*})}{T}\right)$ (3)
in Appendix E, where we assume that $f^{*}\triangleq\inf f(W)>-\infty$. Since diagonal AdaGrad and SignSGD achieve just the same convergence rate in nonconvex settings up to logarithmic factors (Bernstein et al., 2018; Sun et al., 2023; Liu et al., 2024), we may expect that ASGO could achieve a comparable convergence rate with (3) obtained by Muon. It is an interesting future topic to prove this conjecture and further explore the nonconvex behavior of ASGO theoretically.
Remark 3.
Note that a nonconvex smooth convergence analysis of Muon has been presented in Li and Hong (2025). They establish the convergence results under a more general setting, involving gradient noise and momentum following the analysis in Cutkosky and Mehta (2020). However, if we only look at the deterministic case, their result is worse than (3) because of the explicit dependence on dimension $n$. The key here is that the standard smoothness condition with respect to the Frobenius norm is not a good fit for analyzing structured gradient algorithms like Muon or ASGO. Using Assumption 2, we obtain better convergence results for Muon in (3).
A practical implementation of ASGO.
We present a practical implementation of ASGO in Algorithm 2. Our implementation incorporates several common modifications to enhance computational efficiency and stability. (i) Position the preconditioner on the right side, which better aligns with the neuron architecture in DNNs as discussed in Section 5.222The convergence guarantees remain valid since this modification is equivalent to applying the original ASGO with a transposition of the parameter matrix $W$. Notably, right-side preconditioning requires significantly less computational effort than left-side preconditioning, particularly for 1D parameters such as bias terms and layer normalization weights. Left-side preconditioning would correspond to full-matrix AdaGrad in the 1D case, resulting in higher memory consumption and computational complexity. In contrast, right-side preconditioning only requires maintaining a scalar value and implementing AdaGrad-Norm (Streeter and McMahan, 2010; Ward et al., 2020), a scalar step size version of AdaGrad. This approach enables ASGO to efficiently handle all parameter structures within a unified framework, unlike methods such as Muon that require defining separate optimizers (e.g., AdamW (Loshchilov and Hutter, 2017)) for 1D parameters. Although we have not empirically verified whether this AdaGrad-Norm update rule for 1D parameters outperforms AdamW, future work could explore a hybrid approach—similar to Muon—where AdamW is used specifically for 1D parameters while maintaining ASGO for higher-dimensional structures. (ii) Employ exponential moving averages to update the preconditioner and incorporate momentum, following established practice, which results in Adam’s improvements across various tasks over AdaGrad. (iii) Compute and store the inverse of preconditioner matrix every $\tau$ iterations, as is done in distributed Shampoo (Shi et al., 2023), SOAP (Vyas et al., 2024) and Galore (Zhao et al., 2024). This modification increases memory requirements but substantially reduces computational overhead while improving algorithmic stability.
(a) Hessian of Query (4 heads)
(b) Hessian of Key (4 heads)
(c) Hessian of Value (4 heads)
Figure 2: This figure is from Zhang et al. (2024b), describing the Hessian of different parameter blocks in a small Transformer at 1% training step. The near-block-diagonal structure maintains throughout training. But different parameter blocks have different numbers of small dense matrices, where Query and Key correspond to the number of heads.
Special Design for Transformers.
Furthermore, we introduce a specialized adaptation of ASGO for query and key matrices in attention layers. Recent work by Zhang et al. (2024b) has demonstrated that the Hessian structure of query and key layers differs significantly from conventional MLP layers. As illustrated in Figure 2, the number of dense blocks corresponds to the number of attention heads rather than output neurons. This observation aligns with the forward computation of multi-head attention, where attention scores are computed independently across different subspaces, suggesting that query and key parameters could be optimized in a head-wise manner. To leverage this insight, we reshape query and key parameters from matrices in $R^{n\times hd}$ to three-dimensional tensors $R^{h\times n\times d}$, and apply our optimization algorithm independently to each head’s subspace. This restructuring reduces both memory consumption and computational complexity by decreasing the matrix size from $O(h^{2}d^{2})$ into $O(hd^{2})$. For instance, in NanoGPT (Karpathy, 2022), this adjustment reduces the preconditioning size of a single query/key parameter from approximately $10^{6}$ elements into $10^{5}$. The empirical benefits of this modification are evaluated in Section 7.
A Diagonal Variant of ASGO.
Drawing inspiration from Duchi et al. (2011), we also implement a variant of ASGO using diagonal matrices, which we denote as DASGO in Algorithm 3. DASGO can be viewed as a light version of ASGO, which eliminates the need to compute the square root and inverse of matrices and reduces the memory requirements to a comparable level with Adam-mini (Zhang et al., 2024b). However, since DASGO only employs a diagonal preconditioner, it fails to recover the superior theoretical properties of ASGO under block-wise diagonal Hessian settings. We further empirically examine this tradeoff in the next section.
Algorithm 2 A Practical Implementation of ASGO
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, $\epsilon\in\mathbb{R}$, schedule $\{\eta_{t}\}$, momentums $\beta_{1},\beta_{2}\in[0,1)$, and batch size $M\in\mathbb{N}$, update interval $\tau\in\mathbb{N}$ ($\epsilon$ should be small, similar to the $\epsilon$ for Adam or AdaGrad)
2:  Initialize $M_{-1}=0\in\mathbb{R}^{m\times n}$, $V_{-1}=0\in\mathbb{R}^{n\times n}$
3:  for $t=0$ to $T-1$ do
4:     Sample mini-batch $\mathcal{B}_{t}$ with $\left|\mathcal{B}_{t}\right|\equiv M$ uniformly
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$triangleright Compute stochastic gradient
6:     $M_{t}=\beta_{1}M_{t-1}+(1-\beta_{1})G_{t}$triangleright Accumulate momentum
7:     $V_{t}=\beta_{2}V_{t-1}+(1-\beta_{2})G_{t}^{\top}G_{t}$triangleright Use exponential moving average to update $V_{t}$
8:     if $\mathrm{mod}(t,\tau)=0$triangleright Update $\Lambda_{t}^{\mathrm{inv}}$ every $\tau$ iterations
9:      $\Lambda_{t}^{\mathrm{inv}}=\left(V_{t}+\epsilon I_{n}\right)^{-\frac{1}{2}}$
10:     else
11:      $\Lambda_{t}^{\mathrm{inv}}=\Lambda_{t-1}^{\mathrm{inv}}$
12:     end if
13:     $W_{t+1}=W_{t}-\eta_{t}M_{t}\Lambda_{t}^{\mathrm{inv}}$triangleright Precondition from the right side
14:  end for
Algorithm 3 Implementation of DASGO (Diagonal Adaptive Structured Gradient Optimization)
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, $\epsilon\in\mathbb{R}$, schedule $\{\eta_{t}\}$, momentums $\beta_{1},\beta_{2}\in[0,1)$, and batch size $M\in\mathbb{N}$
2:  Initialize $M_{-1}=0\in\mathbb{R}^{m\times n}$, $v_{-1}=0\in\mathbb{R}^{n}$
3:  for $t=0$ to $T-1$ do
4:     Sample mini-batch $\mathcal{B}_{t}$ with $\left|\mathcal{B}_{t}\right|\equiv M$ uniformly
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$triangleright Compute stochastic gradient
6:     $M_{t}=\beta_{1}M_{t-1}+(1-\beta_{1})G_{t}$triangleright Accumulate momentum
7:     $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\operatorname{diag}(G_{t}^{\top}G_{t}$) $\triangleright$ Use exponential moving average to update vector $V_{t}$
8:     $W_{t+1}=W_{t}-\eta_{t}M_{t}\operatorname{diag}(v_{t}+\epsilon)^{-\frac{1}{2}}$triangleright Precondition by a diagonal Matrix from right side
9:  end for
7 Empirical Results
We empirically evaluate the effectiveness of ASGO (Algorithm 2) and DASGO (Algorithm 3) on pretraining tasks for deep learning models. We compare the methods against established optimizers, including AdamW (Kingma and Ba, 2014; Loshchilov and Hutter, 2017), Shampoo (Gupta et al., 2018), and Muon (Jordan et al., 2024). Several important implementation details should be noted for fair comparison. First, since Muon is designed to operate exclusively on matrix and tensor parameters, we follow Jordan et al. (2024) to apply AdamW update rules to all 1D parameters within the Muon optimizer to ensure that it can handle the complete model. Additionally, the specialized attention layer modification for query and key matrices described in Section 6 can be applied to Muon as well, providing similar benefits in reducing memory consumption and computational complexity.
Recent research (Shi et al., 2023; Morwani et al., 2024; Lin et al., 2024) has suggested treating the inverse order in Shampoo as a tunable hyperparameter rather than using the standard value of $1/4$. Notably, Morwani et al. (2024) demonstrated that Shampoo2 (Shampoo with inverse-order $1/2$) provides a superior approximation of the full-matrix AdaGrad optimizer. Motivated by this finding, we treat the inverse order as a hyperparameter in our pretraining experiments. Additionally, Shampoo’s performance is highly dependent on the initialization of its preconditioner matrices, as a well-initialized preconditioner is crucial for accurately approximating the Empirical Fisher Information Matrix (Morwani et al., 2024). To address this initialization sensitivity, we implement a preconditioner warmup phase specifically for the Shampoo optimizer. Following the methodology in Ren and Goldfarb (2021), we dedicate the first epoch exclusively to accumulating statistics for the preconditioner matrices without updating model parameters. This approach yields a more robust estimation of the preconditioner, which significantly improves the stability of the Shampoo algorithm during subsequent training iterations.
For consistency across all optimization methods, we employ the OneCycleLR learning rate schedule, which has been shown to provide stable convergence properties in deep learning tasks.
7.1 Pretraining NanoGPT
We conducted experiments using the NanoGPT architecture on the Shakespeare character dataset, following the configuration in (Karpathy, 2022). To ensure a fair comparison, we maintained consistent hyperparameter (HP) settings across all optimization methods, varying only those parameters that significantly impacted optimizer performance, i.e., learning rate, 1st and 2nd order moment coefficients ($\beta_{1}$ and $\beta_{2}$), preconditioner update frequency ($\tau$), and warm-up ratio  [^3] A full description of how these “best” HPs were selected and their values is presented in Appendix A.. Figure 3 presents a comparison of the pretraining performance of ASGO and DASGO with that of Shampoo, Muon and AdamW on the NanoGPT model. Examining both train and test loss curves, ASGO consistently outperforms Shampoo despite requiring only half the memory consumption and computational complexity, which highlights ASGO’s practical advantages for training language models. Furthermore, ASGO and Muon achieve the lowest final training and test losses, outperforming all other methods. This consistent performance between ASGO and Muon aligns well with our discussions in Section 6.
The lightweight DASGO optimizer achieves competitive results against AdamW in both training and test loss metrics. DASGO demonstrates particularly strong performance during the initial training phase (first 200 training step), where it converges faster than other methods. This early convergence advantage suggests that DASGO could serve as an effective warmup optimizer in multi-stage training pipelines. On the other hand, ASGO demonstrates more stable convergence with less oscillation in later training stages compared to DASGO. This performance gap between DASGO and ASGO highlights the value of preserving non-diagonal elements in the preconditioner matrices, particularly for capturing parameter interactions in attention-based architectures. Nevertheless, DASGO’s competitive performance while requiring substantially fewer computational resources makes it an attractive option for resource-constrained scenarios.
(a) Train Loss
(b) Test Loss
Figure 3: Train Loss and Test Loss on the NanoGPT and Shakespeare Character Dataset
Ablation Study: Impact of Special Query/Key Design
In Section 6, we highlighted the computational benefits of specialized processing for Query and Key matrices in Transformer architectures. To validate these theoretical advantages empirically, we conducted an ablation study comparing the performance and training stability of ASGO and Muon, both with and without the specialized Query-Key processing. We maintained the optimal hyperparameter configuration established in Section 7.1 for all parameters except the learning rate. For each algorithm variant (with and without specialized processing), we randomly sampled learning rates from a log-uniform distribution ranging from $10^{-5}$ to $10^{-1}$. Each configuration was trained for 5 epochs on the NanoGPT model, after which we recorded the validation loss. This approach allowed us to evaluate both performance and robustness to learning rate selection. Figure 4 presents the distribution of validation losses after 5 epochs for both algorithms. The violin plots [^4] Violin plot outlines depict empirical probability density; i.e., the width of the shaded area represents the proportion of the data located there. Box plots within a violin plot display the median and inter-quartile range. reveal several key insights: (1) For ASGO (Figure 4a), the benefits of specialized Query-Key processing are substantial. The specialized implementation demonstrates a narrower, lower distribution of validation losses (centered around 2.5), indicating greater stability across different learning rates. In contrast, the vanilla implementation shows a long-tailed distribution extending beyond loss values of 6.0, with several outliers representing training instability at certain learning rates. (2) For Muon (Figure 4b), both implementations demonstrate nearly identical performance distributions, with their median validation losses and distribution shapes showing minimal differences. However, the specialized implementation achieves this comparable performance while reducing memory requirements and computational complexity. This represents a clear efficiency advantage—the specialized Query-Key processing effectively provides computational savings with no performance penalty.
(a) ASGO
(b) Muon
Figure 4: Distribution of validation losses after 5 epochs with varying learning rates.
8 Conclusions
In this paper, we propose a novel algorithm ASGO, aiming to utilize the structured properties of optimization problems. We prove that ASGO achieves significantly better convergence rates compared to full-matrix AdaGrad and Shampoo. Based on the theory, we demonstrate that ASGO can benefit from low-rank gradients and block-wise diagonal Hessians, which are widely observed structured properties of deep neural networks. We further propose some practical modifications on ASGO, and empirical results on language model tasks verify the effectiveness of our approach. For future work, we plan to extend our research in two key directions:
  • Larger Foundation Models: We aim to evaluate the algorithm’s performance on large foundation models to demonstrate its effectiveness in pretraining LLMs.
  • We intend to generalize our algorithm to handle tensor cases, which would significantly expand its applicability across a broader range of neural network architectures.
Appendix A Details of Empirical Experiments
For our experiments in Section 7.1, we employ the NanoGPT architecture as given in Karpathy [2022]. The model consists of 6 Transformer layers, 6 attention heads, and an embedding dimension of 384. We utilize the OneCycle learning rate schedule consistently across all optimization algorithms to ensure fair comparison. Training is conducted on the Shakespeare character-level dataset with a sequence length of 256 tokens.
To ensure optimal performance for each optimizer, we conduct a comprehensive hyperparameter search using random search  [Bergstra and Bengio, 2012, Choi et al., 2019] within the following predefined ranges: learning rate(lr): $[10^{-5},10^{-1}]$; $\beta_{1}$: $[0.7,0.99]$; $\beta_{2}$: $[0.7,0.99]$; and select the warmup factor from the set $[0.1,0.15,0.2,0.25,0.3]$. For the optimizers Shampoo and ASGO, we select the update frequency $\tau$ from the list $[5,10,15]$. We also treat the inverse order in Shampoo as a hyperparameter and select it as either $-\frac{1}{2}$ or $-\frac{1}{4}$. HPs were selected based on validation loss after 20 epochs of training. For each optimizer, we performed a random search for 16 hours to find the optimal HP combination. These best HP choices for training NanoGPT are given in Table 1.
Optimizer learning rate $\beta_{1}$ $\beta_{2}$ warmup factor update frequency inverse order
Muon 0.00349 0.9881 N/A 0.3 N/A N/A
AdamW 0.00450 0.9332 0.9528 0.2 N/A N/A
DASGO 0.060 0.9584 0.9435 0.2 N/A N/A
Shampoo 0.00593 0.9402 0.9760 0.3 15 $-\frac{1}{4}$
ASGO 0.01470 0.9541 0.8487 0.3 15 N/A
Table 1: Hyperparameter selection for NanoGPT experiment
Appendix B Auxiliary Lemmas
Lemma 1 (Trace properties).
For arbitrary matrices $A\in\mathbb{R}^{m\times n}$, $B\in\mathbb{R}^{n\times m}$, $X,Y\in\mathbb{R}^{m\times m}$, we have the following basic properties for the trace:
  1. 1.
    ${\rm tr}\left(X\right)={\rm tr}\left(X^{\top}\right)$;
  2. 2.
    ${\rm tr}\left(AB\right)={\rm tr}\left(BA\right)$;
  3. 3.
    if $X$ is symmetric positive semidefinite, ${\rm tr}\left(X\right)=\left\|X\right\|_{*}\geq 0$;
  4. 4.
    if $X,Y\succeq 0$ and $X$ is symmetric, ${\rm tr}\left(XY\right)\geq 0$.
These are standard trace properties. The following lemma notes the operator monotonicity of the power functions, which is a classic result [Löwner, 1934, Ando et al., 2004, Gupta et al., 2018].
Lemma 2.
The function $f:x\to x^{\alpha}$ with $\alpha\in[0,1]$ is operator-monotone, i.e. if $0\preceq A\preceq B$, it holds that $A^{\alpha}\preceq B^{\alpha}$.
Lemma 3.
For symmetric positive definite matrices $X,Y\in\mathbb{R}^{m\times m}$, it holds that
${\rm tr}\left((X+Y)^{\frac{1}{2}}\right)\leq{\rm tr}\left(X^{\frac{1}{2}}+Y^{\frac{1}{2}}\right).$
Proof.
It holds that
${\rm tr}\left((X+Y)^{\frac{1}{2}}\right)=$ ${\rm tr}\left((X+Y)^{-\frac{1}{2}}(X+Y)\right)$
$=$ ${\rm tr}\left((X+Y)^{-\frac{1}{2}}X\right)+{\rm tr}\left((X+Y)^{-\frac{1}{2}}Y\right)$
$\geq$ ${\rm tr}\left(X^{\frac{1}{2}}\right)+{\rm tr}\left(Y^{\frac{1}{2}}\right),$
where the last inequality is based on Lemma 2 such that $(X+Y)^{\frac{1}{2}}\succeq X^{\frac{1}{2}}$ and the fact that $A^{-1}\preceq B^{-1}$ if $A\succeq B$ for symmetric positive definite matrices $A,B$. Further based on Lemma 1, we can finish the proof. ∎
Lemma 4.
For a symmetric positive semidefinite matrix $X\in\mathbb{R}^{m\times m}$, it holds that
${\rm tr}\left(X^{\frac{1}{2}}\right)\leq\sum_{j=1}^{m}\sqrt{[X]_{j,j}}.$
Proof.
The inequality is equivalent to that
$\sum_{i=1}^{m}\sqrt{\lambda_{i}(X)}\leq\sum_{i=1}^{m}\sqrt{[X]_{i,i}},$
where $\lambda_{i}(X)$ denotes the $i$-th largest eigenvalue of $X$ (the same as singular values for real symmetric positive semidefinite matrices). Firstly, we have the fact that the eigenvalues $\{\lambda_{i}(X)\}_{i=1}^{m}$ majorize the diagonal entries $\{[X]_{i,i}\}_{i=1}^{m}$, i.e. for all $1\leq l<m$,
$\sum_{i=1}^{l}\lambda_{i}(X)\geq\sum_{i=1}^{l}[X]_{i,i},\quad\text{and}\quad\sum_{i=1}^{m}\lambda_{i}(X)=\sum_{i=1}^{m}[X]_{i,i}.$
Also, we know $g:\{x_{i}\}_{i=1}^{m}\to\sum_{i=1}^{m}x_{i}^{\frac{1}{2}}$ is a Schur-concave operator. Then we obtain the equality based on the Schur-Horn theorem, which implies that for a Schur-concave operator $g$, if sequence $\{x_{i}\}_{i=1}^{m}$ majorizes $\{y_{i}\}_{i=1}^{m}$, then $g(\{x_{i}\}_{i=1}^{m})\leq g(\{y_{i}\}_{i=1}^{m})$. ∎
Lemma 5.
$\left\|\cdot\right\|_{L}$ is a norm, i.e., it satisfies the basic properties of norms. Its dual norm is $\left\|\cdot\right\|_{L^{-1}}$.
Proof.
Denote $X=[x_{1},\dots,x_{n}]\in\mathbb{R}^{m\times n}$, where each $x_{i}\in\mathbb{R}^{1\times m}$. Then we have
${\rm tr}\left(X^{\top}LX\right)=\sum_{i=1}^{n}x_{i}^{\top}Lx_{i}=\begin{bmatrix}x_{1}\\ \cdots\\ x_{n}\end{bmatrix}^{\top}\begin{bmatrix}L&&&\\ &L&&\\ &&\cdots&\\ &&&L\end{bmatrix}\begin{bmatrix}x_{1}\\ \cdots\\ x_{n},\end{bmatrix},$
which is a norm for the space $\mathbb{R}^{m\times n}$ whose dual norm is $\left\|\cdot\right\|_{L^{-1}}$. This concludes the proof. ∎
Lemma 6.
Assume a non-negative sequence $\{x_{j}\}_{j=1}^{n}$ and a positive sequence $\{s_{j}\}_{j=1}^{n}$ with $S=\sum_{i=1}^{n}s_{j}$, it holds that
$\frac{1}{S}\sum_{j=1}^{n}x_{j}\leq\sqrt{\frac{1}{S}\sum_{j=1}^{n}\frac{x_{j}^{2}}{s_{j}}}.$ (4)
The inequality holds as an equality if and only if for all $i=1,\cdots,n$ and $j=1,\cdots,n$,
$\frac{x_{i}}{s_{i}}=\frac{x_{j}}{s_{j}}.$
Proof.
A proof can be found in the ICLR version of Liu et al. [2024] for Lemma G.5. ∎
Appendix C Proof of Nonsmooth Convergence of ASGO
Proof of Theorem 1.
Denote $\Delta W_{t}\triangleq W_{t}-W_{*}$ where $W_{*}$ is an optimum of Problem (1). From the algorithm update, we have
$\Delta W_{t+1}^{\top}\Lambda_{t}\Delta W_{t+1}=\Delta W_{t}^{\top}\Lambda_{t}\Delta W_{t}-\eta_{t}\left(G_{t}^{\top}\Delta W_{t}+\Delta W_{t}^{\top}G_{t}\right)+\eta_{t}^{2}G_{t}^{\top}\Lambda_{t}^{-1}G_{t}.$
Then, by taking trace of the above equality and rearranging, we can obtain that
$2\eta_{t}{\rm tr}\left(G_{t}^{\top}\Delta W_{t}\right)={\rm tr}\left(\Delta W_{t}^{\top}\Lambda_{t}\Delta W_{t}-\Delta W_{t+1}^{\top}\Lambda_{t}\Delta W_{t+1}\right)+\eta_{t}^{2}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right).$
By convexity, we know
$\mathbb{E}\left[{\rm tr}\left(G_{t}^{\top}\Delta W_{t}\right)\right]=\mathbb{E}\left[\left\langle\nabla f(W_{t}),\Delta W_{t}\right\rangle\right]\geq\mathbb{E}[f(W_{t})]-f(W_{*}).$
Then combining these and taking summation over $t$ and taking $\eta_{t}\equiv\eta$, we have
$\begin{split}&2\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})]-f(W_{*})\\ \leq&\frac{1}{\eta}\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}^{\top}\Lambda_{t}\Delta W_{t}-\Delta W_{t+1}^{\top}\Lambda_{t}\Delta W_{t+1}\right)\right]+\eta\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)\right]\\ =&\frac{1}{\eta}\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}^{\top}\Lambda_{t}\Delta W_{t}-\Delta W_{t}^{\top}\Lambda_{t-1}\Delta W_{t}\right)\right]+\frac{1}{\eta}{\rm tr}\left(\Delta W_{0}^{\top}\Lambda_{-1}\Delta W_{0}\right)-\frac{1}{\eta}\mathbb{E}\left[{\rm tr}\left(\Delta W_{T}^{\top}\Lambda_{T-1}\Delta W_{T}\right)\right]\\ &+\eta\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)\right]\\ \leq&\frac{1}{\eta}{\rm tr}\left(\Delta W_{0}^{\top}\Lambda_{-1}\Delta W_{0}\right)+\frac{1}{\eta}\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}^{\top}(\Lambda_{t}-\Lambda_{t-1})\Delta W_{t}\right)\right]+\eta\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)\right],\end{split}$ (5)
where we note $\Lambda_{-1}=\epsilon I_{m}$. Thus the first term on the RHS of (5) can be bounded by $\epsilon$ as
$\frac{1}{\eta}{\rm tr}\left(\Delta W_{0}^{\top}\Lambda_{-1}\Delta W_{0}\right)=\frac{\epsilon}{\eta}\left\|\Delta W_{0}\right\|_{\mathrm{F}}^{2}\leq\frac{\epsilon D_{\mathrm{F}}^{2}}{\eta}.$ (6)
We then deal with the second and third terms separately. For the second term, we have
$\begin{split}\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}^{\top}(\Lambda_{t}-\Lambda_{t-1})\Delta W_{t}\right)=&\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}\Delta W_{t}^{\top}(\Lambda_{t}-\Lambda_{t-1})\right)\\ =&\sum_{t=0}^{T-1}\left\langle\Delta W_{t}\Delta W_{t}^{\top},\Lambda_{t}-\Lambda_{t-1}\right\rangle\\ \leq&\sum_{t=0}^{T-1}\left\|\Delta W_{t}\Delta W_{t}^{\top}\right\|_{\mathrm{op}}\left\|\Lambda_{t}-\Lambda_{t-1}\right\|_{*}\\ =&\sum_{t=0}^{T-1}\left\|\Delta W_{t}\right\|_{\mathrm{op}}^{2}{\rm tr}\left(\Lambda_{t}-\Lambda_{t-1}\right)\leq D_{\mathrm{op}}^{2}{\rm tr}\left(\Lambda_{T-1}-{\rm tr}\left(\Lambda_{-1}\right)\right).\end{split}$ (7)
Note that the first and last equalities are based on the properties of the trace, (see Lemma 1). The first inequality is based on the duality of the $\left\|\cdot\right\|_{\mathrm{op}}$ and $\left\|\cdot\right\|_{*}$ norms. The third equality relies on the positive semidefiniteness of $\Lambda_{t}-\Lambda_{t-1}$ for all $t$ based on Lemma 2, since $\Lambda_{t}^{2}-\Lambda_{t-1}^{2}=G_{t}G_{t}^{\top}\succeq 0$.
For the third term of (5), we have
$\begin{split}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)=&{\rm tr}\left(G_{t}G_{t}^{\top}\Lambda_{t}^{-1}\right)={\rm tr}\left((\Lambda_{t}^{2}-\Lambda_{t-1}^{2})\Lambda_{t}^{-1}\right)\\ =&{\rm tr}\left(\left((\Lambda_{t}-\Lambda_{t-1})\cdot 2\Lambda_{t}-(\Lambda_{t}-\Lambda_{t-1})^{2}+\Lambda_{t-1}\Lambda_{t}-\Lambda_{t}\Lambda_{t-1}\right)\Lambda_{t}^{-1}\right)\\ =&2{\rm tr}\left(\Lambda_{t}-\Lambda_{t-1}\right)-{\rm tr}\left((\Lambda_{t}-\Lambda_{t-1})^{2}\Lambda_{t}^{-1}\right)+{\rm tr}\left((\Lambda_{t-1}\Lambda_{t}-\Lambda_{t}\Lambda_{t-1})\Lambda_{t}^{-1}\right)\\ \leq&2{\rm tr}\left(\Lambda_{t}-\Lambda_{t-1}\right).\end{split}$ (8)
where the last inequality follows from the fact that
${\rm tr}\left((\Lambda_{t}-\Lambda_{t-1})^{2}\Lambda_{t}^{-1}\right)={\rm tr}\left(\left[(\Lambda_{t}-\Lambda_{t-1})\Lambda_{t}^{-\frac{1}{2}}\right]\left[(\Lambda_{t}-\Lambda_{t-1})\Lambda_{t}^{-\frac{1}{2}}\right]^{\top}\right)\geq 0$
and
${\rm tr}\left((\Lambda_{t-1}\Lambda_{t}-\Lambda_{t}\Lambda_{t-1})\Lambda_{t}^{-1}\right)={\rm tr}\left(\Lambda_{t-1}\right)-{\rm tr}\left(\Lambda_{t}\Lambda_{t-1}\Lambda_{t}^{-1}\right)=0,$
which are based on the positive definiteness of $\Lambda_{t}$ and the properties of the trace (Lemma 1). Therefore, by substituting (6), (7), and (8) into (5), we have
$2\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})]-f(W_{*})$
$\leq$ $\frac{1}{\eta}{\rm tr}\left(\Delta W_{0}^{\top}\Lambda_{-1}\Delta W_{0}\right)+\frac{1}{\eta}\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}^{\top}(\Lambda_{t}-\Lambda_{t-1})\Delta W_{t}\right)\right]+\eta\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)\right]$
$\leq$ $\frac{\epsilon D_{\mathrm{F}}^{2}}{\eta}+\left(\frac{D_{\mathrm{op}}^{2}}{\eta}+2\eta\right)\mathbb{E}\left[{\rm tr}\left(\Lambda_{T-1}-\Lambda_{-1}\right)\right]$
$=$ $\left(\frac{D_{\mathrm{op}}^{2}}{\eta}+2\eta\right)\mathbb{E}\left[{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)\right]+\frac{\epsilon D_{\mathrm{F}}^{2}}{\eta}.$
Taking $\eta=D_{\mathrm{op}}$ completes the proof.
Based on Theorem 1 and Lemma 8, we can also prove Corollary 2.
Proof of Corollary 2.
Based on the results in Theorem 1, if we additionally have
$\mathbb{E}\left[G_{t}G_{t}^{\top}\right]\preceq Q^{2},$
then using Lemma 8, we can obtain that
[…]
$=$ $\mathbb{E}\left[\sqrt{\left\|Q\right\|_{*}{\rm tr}\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}Q^{-1}\right)}\right]$
$\leq$ $\sqrt{\left\|Q\right\|_{*}\sum_{t=0}^{T-1}{\rm tr}\left(\mathbb{E}\left[G_{t}G_{t}^{\top}\right]Q^{-1}\right)}$
$\leq$ $\sqrt{\left\|Q\right\|_{*}\sum_{t=0}^{T-1}\left\|Q\right\|_{*}}=\sqrt{T}\left\|Q\right\|_{*},$
where the first equality is based on Lemma 1 and the second inequality is based on the fact that $g(x)=\sqrt{x}$ is concave for $x\geq 0$. This concludes the proof.
We also involve the following proof for the comparison between Theorem 1 and the convergence rates of Shampoo and full-matrix AdaGrad.
Proof of Comparison with Shampoo and full-matrix AdaGrad.
We list the convergence rates here.
Full-Matrix AdaGrad: $\quad\mathcal{O}\left(D_{\mathrm{F}}\sum_{j=1}^{m}\sum_{i=1}^{n}\sqrt{\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right)$
Shampoo: $\quad\mathcal{O}\left(\sqrt{r_{G}}D_{\mathrm{F}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)^{\frac{1}{4}}\right)\right)$
ASGO: $\quad\mathcal{O}\left(D_{\mathrm{op}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)\right)\leq\mathcal{O}\left(D_{\mathrm{op}}\cdot\sum_{j=1}^{m}\sqrt{\sum_{i=1}^{n}\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right)$
The last inequality for ASGO is based on Lemma 4. We first compare ASGO and full-matrix AdaGrad:
${\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)\leq\sum_{j=1}^{m}\sqrt{\left[\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right]_{j,j}}=\sum_{j=1}^{m}\sqrt{\sum_{i=1}^{n}\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\leq\sum_{j=1}^{m}\sum_{i=1}^{n}\sqrt{\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}},$
where the first inequality is based on Lemma 4 and the second inequality is simply a fact that for a vector $x$, we have $\left\|x\right\|_{2}\geq\left\|x\right\|_{1}$. Thus we prove that the proven convergence rate of ASGO is at least $D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than full-matrix AdaGrad.
Then we compare ASGO and Shampoo. We have
${\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)=$ $\left\langle\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}},\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right\rangle$
$\leq$ ${\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot\left\|\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right\|_{\mathrm{op}}$
$=$ ${\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot\left\|\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right\|_{\mathrm{op}}^{\frac{1}{4}},$
where the inequality is based on the fact that $\left\|\cdot\right\|_{\mathrm{op}}$ and $\left\|\cdot\right\|_{*}$ are dual norms. Then we have
$\left\|\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right\|_{\mathrm{op}}^{\frac{1}{4}}\leq$ $\left(\sum_{t=0}^{T-1}\left\|G_{t}G_{t}^{\top}\right\|_{\mathrm{op}}\right)^{\frac{1}{4}}=\left(\sum_{t=0}^{T-1}\left\|G_{t}^{\top}G_{t}\right\|_{\mathrm{op}}\right)^{\frac{1}{4}}$
$\leq$ $\left({\rm tr}\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)\right)^{\frac{1}{4}}\leq{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)^{{}^{\frac{1}{4}}}\right),$
where the first inequality is based on that $\left\|\cdot\right\|_{\mathrm{op}}$ is a norm and the second inequality is based on the fact that ${\rm tr}\left(X\right)\geq\left\|X\right\|_{\mathrm{op}}$ for symmetric positive semidefinite $X$, and the third inequality is based on the fact that $({\rm tr}\left(X\right))^{\frac{1}{4}}\leq{\rm tr}\left(X^{\frac{1}{4}}\right)$ because if we denote $\sigma_{j}$ as the $j$-th singular value of $X$, we have
$({\rm tr}\left(X\right))^{\frac{1}{4}}=\left(\sum_{j=1}^{r}\sigma_{j}\right)^{\frac{1}{4}}\leq\sum_{j=1}^{r}\sigma_{j}^{\frac{1}{4}}.$
Thus we can conclude that
${\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)\leq{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)^{\frac{1}{4}}\right).$
Therefore, the proven rate of ASGO is at least $\sqrt{r_{G}}D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than Shampoo.
Appendix D Proof of Smooth Convergence of ASGO
The starting point of the smooth analysis is Theorem 1. For notation simplicity, we denote $N_{t}\triangleq G_{t}-\nabla f(W_{t})$ and $\nabla f_{t}\triangleq\nabla f(W_{t})$ in this section.
Lemma 7 (Separate Gradient and Noise).
Under the same settings as Theorem 1, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]\leq$ $\frac{D_{\mathrm{op}}}{T}\mathbb{E}\left[\left\|\left(2\sum_{t=0}^{T-1}\nabla f_{t}\nabla f_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]+\frac{D_{\mathrm{op}}}{T}\mathbb{E}\left[\left\|\left(2\sum_{t=0}^{T-1}N_{t}N_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}.$
Proof.
Based on Theorem 1, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]\leq$ $\frac{D_{\mathrm{op}}}{T}\mathbb{E}\left[\left\|\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}$
$\leq$ $\frac{D_{\mathrm{op}}}{T}\mathbb{E}\left[\left\|\left(\sum_{t=0}^{T-1}2\nabla f_{t}\nabla f_{t}^{\top}+2N_{t}N_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}$
$\leq$ $\frac{D_{\mathrm{op}}}{T}\mathbb{E}\left[\left\|\left(2\sum_{t=0}^{T-1}\nabla f_{t}\nabla f_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]+\frac{D_{\mathrm{op}}}{T}\mathbb{E}\left[\left\|\left(2\sum_{t=0}^{T-1}N_{t}N_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}.$
Here the first inequality comes directly from Theorem 1. The second inequality is based on the fact that for all $A,B\in\mathbb{R}^{m\times n}$ we have
$2AA^{\top}+2BB^{\top}-(A+B)(A+B)^{\top}=$ $AA^{\top}+BB^{\top}-A^{\top}B-B^{\top}A$
$=$ $(A-B)(A-B)^{\top}\succeq 0.$
The last inequality is based on Lemma 3. ∎
The following lemma is a key technical lemma for proving the smooth convergence results, which is a generalization to the matrix case of Lemma 6.
Lemma 8 (An upper bound on $\left\|\cdot\right\|_{*}$).
For a symmetric positive definite matrix $\Lambda\in\mathbb{R}^{m\times m}$ and matrix $G\in\mathbb{R}^{m\times n}$, it holds that
$\left\|G\right\|_{*}\leq\sqrt{\left\|\Lambda\right\|_{*}{\rm tr}\left(G^{\top}\Lambda^{-1}G\right)}.$ (9)
Proof.
We first consider the case $\Lambda=\text{diag}[\lambda_{1},\dots,\lambda_{m}]$ is a diagonal matrix. We have
$\left\|G\right\|_{*}=$ ${\rm tr}\left(\left(GG^{\top}\right)^{\frac{1}{2}}\right)\leq\sum_{j=1}^{m}\sqrt{\left[GG^{\top}\right]_{j,j}}$
$\leq$ $\sqrt{\left(\sum_{j=1}^{m}\lambda_{j}\right)\left(\sum_{j=1}^{m}\frac{\left[GG^{\top}\right]_{j,j}}{\lambda_{j}}\right)}$
$=$ $\sqrt{\left\|\Lambda\right\|_{*}{\rm tr}\left(GG^{\top}\Lambda^{-1}\right)}=\sqrt{\left\|\Lambda\right\|_{*}{\rm tr}\left(G^{\top}\Lambda^{-1}G\right)},$
where the first inequality is based on Lemma 4 and the second inequality is based on Lemma 6.
Then we prove that this holds for general symmetric positive definite matrix $\Lambda$. Let the singular value decomposition be $\Lambda=U\Sigma U^{\top}$. Denote $\tilde{G}=U^{\top}G$. Since $\Sigma\in\mathbb{R}^{m\times m}$ is a diagonal matrix, it holds that
$\left\|G\right\|_{*}=$ $\left\|\tilde{G}\right\|_{*}\leq\sqrt{\left\|\Sigma\right\|_{*}{\rm tr}\left(\tilde{G}^{\top}\Sigma^{-1}\tilde{G}\right)}=\sqrt{\left\|\Lambda\right\|_{*}{\rm tr}\left(G^{\top}U\Sigma^{-1}U^{\top}G\right)}$
$=$ $\sqrt{\left\|\Lambda\right\|_{*}{\rm tr}\left(G^{\top}\Lambda^{-1}G\right)},$
which concludes the proof. ∎
We also use the following lemma to indicate the reduction of variance by batch size $M$.
Lemma 9 (Variance reduction by batch size).
Under Assumption 3, we have
$\mathbb{E}\left[N_{t}N_{t}^{\top}\right]\preceq\frac{1}{M}V^{2},$
where we denote $N_{t}\triangleq G_{t}-\nabla f(W_{t})$ for Algorithm 1.
Proof.
For notation simplicity, we denote $N(W_{t};\xi)\triangleq\nabla f(W_{t};\xi)-\nabla f(W_{t})$ and thus
$N_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}}N(W_{t};\xi).$
Then it holds that
$\mathbb{E}\left[N_{t}N_{t}^{\top}\right]=$ $\frac{1}{M^{2}}\mathbb{E}\left[\left(\sum_{\xi\in\mathcal{B}}N(W_{t};\xi)\right)\left(\sum_{\zeta\in\mathcal{B}}N(W_{t};\zeta)\right)^{\top}\right]$
$=$ $\frac{1}{M^{2}}\sum_{\xi\in\mathcal{B}}\sum_{\zeta\in\mathcal{B}}\mathbb{E}\left[N(W_{t};\xi)N(W_{t};\zeta)^{\top}\right]$
$=$ $\frac{1}{M^{2}}\sum_{\xi\in\mathcal{B}}\mathbb{E}\left[N(W_{t};\xi)N(W_{t};\xi)^{\top}\right]$
$\preceq$ $\frac{1}{M^{2}}\sum_{\xi\in\mathcal{B}}V^{2}=\frac{1}{M}V^{2},$
where the last equality is based on that $\nabla f(W_{t};\xi)$ are mutually independent and $\mathbb{E}[N(W_{t};\xi)]=0$ and the last inequality is based on Assumption 3. ∎
Then based on these lemmas, we can prove the smooth results.
Proof of Theorem 3.
We separately deal with the bias and variance terms, which refer to the first two terms on the RHS of Lemma 7. For the bias part, by plugging in the smoothness matrix $L$ into Lemma 8, we can obtain
[…]
$=$ $\mathbb{E}\left[\sqrt{\left\|L\right\|_{*}{\rm tr}\left(\sum_{t=0}^{T-1}\nabla f_{t}\nabla f_{t}^{\top}L^{-1}\right)}\right]$
$\leq$ $\sqrt{\left\|L\right\|_{*}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\|\nabla f_{t}\right\|_{L^{-1}}^{2}\right]}\leq\sqrt{\left\|L\right\|_{*}\sum_{t=0}^{T-1}2(f(W_{t})-f(W_{*}))},$
where the equalities are based on Lemma 1 and the second inequality is based on the fact that $g(x)=\sqrt{x}$ is concave for $x\geq 0$. The last inequality is based on the properties of smoothness [Nesterov et al., 2018] and the fact that $\left\|\cdot\right\|_{L^{-1}}$ is the dual norm of $\left\|\cdot\right\|_{L}$.
Then, for the variance part, let us first assume $M=1$ for simplicity, then we have
[…]
$=$ $\mathbb{E}\left[\sqrt{\left\|V\right\|_{*}{\rm tr}\left(\sum_{t=0}^{T-1}N_{t}N_{t}^{\top}V^{-1}\right)}\right]$
$\leq$ $\sqrt{\left\|V\right\|_{*}\sum_{t=0}^{T-1}{\rm tr}\left(\mathbb{E}\left[N_{t}N_{t}^{\top}\right]V^{-1}\right)}$
$\leq$ $\sqrt{\left\|V\right\|_{*}\sum_{t=0}^{T-1}\left\|V\right\|_{*}}=\sqrt{T}\left\|V\right\|_{*},$
where the equality is based on Lemma 1. and the second inequality is based on the fact that $g(x)=\sqrt{x}$ is concave for $x\geq 0$. Then plugging these in Lemma 7, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]\leq$ $\frac{D_{\mathrm{op}}}{T}\mathbb{E}\left[\left\|\left(2\sum_{t=0}^{T-1}\nabla f_{t}\nabla f_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]+\frac{D_{\mathrm{op}}}{T}\mathbb{E}\left[\left\|\left(2\sum_{t=0}^{T-1}N_{t}N_{t}^{\top}\right)^{\frac{1}{2}}\right\|_{*}\right]+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}$
$\leq$ $\frac{2D_{\mathrm{op}}}{T}\sqrt{\left\|L\right\|_{*}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]}+\frac{\sqrt{2}D_{\mathrm{op}}\left\|V\right\|_{*}}{\sqrt{T}}+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}$
Thus if we denote $x=\sqrt{\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]}$, we can write the inequality as
$x^{2}\leq bx+c,$
where
$b=\frac{2D_{\mathrm{op}}\sqrt{\left\|L\right\|_{*}}}{\sqrt{T}},\quad c=\frac{\sqrt{2}D_{\mathrm{op}}\left\|V\right\|_{*}}{\sqrt{T}}+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}.$
Then as $x\geq 0$, we can solve this simple quadratic inequality to obtain that
$x^{2}\leq\frac{1}{4}\left(b+\sqrt{b^{2}+4c}\right)^{2}\leq 2b^{2}+2c$
$\Longleftrightarrow\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]\leq\frac{4D_{\mathrm{op}}^{2}\left\|L\right\|_{*}}{T}+\frac{2\sqrt{2}D_{\mathrm{op}}\left\|V\right\|_{*}}{\sqrt{T}}+\frac{2\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T},$
which concludes the proof for $M=1$. Then, based on Lemma 9 to incorporate batch size $M>1$, we finish the proof. ∎
Appendix E Proof of Section 6
Muon [Jordan et al., 2024] is approximately the following algorithm:
$\begin{split}B_{t}=&\mu B_{t-1}+G_{t}\\ U_{t},S_{t},V_{t}=&\text{SVD}(B_{t})\\ W_{t+1}=&W_{t}-\eta_{t}U_{t}V_{t}^{\top}\end{split}$ (10)
where $G_{t}$ is the gradient obtained at $W_{t}$ and $U_{t}\in\mathbb{R}^{m\times r_{t}},V_{t}\in\mathbb{R}^{r_{t}\times n}$ are the left and right singular vectors of $B_{t}$ with $r_{t}$ being the rank of $B_{t}$. Note that Muon implements this algorithm using the Newton-Schulz matrix iteration to approximately compute $U_{t}V_{t}^{\top}$ instead of directly employing SVD to improve computational efficiency. The relation between two original Muon forms (10) and (2) is interpreted in Bernstein and Newhouse [2024] when mentioning spectral descent. In the following discussion, we mention Muon as described in (10). We first address the equivalence of ASGO and Muon without gradient accumulation and momentum.
Proposition 1.
If we set ASGO (Algorithm 2 version) with $\beta_{1}=\beta_{2}=0$ and $\epsilon\to 0$, $\tau=1$ and Muon (10) with $\mu=0$, ASGO is equivalent to Muon.
Proof.
For ASGO with $\beta_{1}=\beta_{2}=0$, the update is
$W_{t+1}=$ $W_{t}-\eta_{t}(G_{t}G_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}G_{t}$
$=$ $W_{t}-\eta_{t}(\tilde{U}_{t}\tilde{S}_{t}\tilde{S}_{t}^{\top}\tilde{U}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{U}_{t}\tilde{S}_{t}\tilde{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}\tilde{U}_{t}(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{U}_{t}^{\top}\tilde{U}_{t}\tilde{S}_{t}\tilde{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}\tilde{U}_{t}(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{U}_{t}^{\top}\tilde{U}_{t}\tilde{S}_{t}\tilde{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}\tilde{U}_{t}(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{S}_{t}\tilde{V}_{t}^{\top}$
where $\tilde{U}_{t}\in\mathbb{R}^{m\times m},\tilde{S}_{t}\in\mathbb{R}^{m\times n},\tilde{V}_{t}\in\mathbb{R}^{n\times n}$ is the full-size SVD of $G_{t}$ such that $G_{t}=\tilde{U}_{t}\tilde{S}_{t}\tilde{V}_{t}^{\top}$. We have $\tilde{S}_{t}\tilde{S}_{t}^{\top}$ is a diagonal matrix with diagonal entries being the square of singular values of $G_{t}$ and zeros. Thus we have each possibly nonzero entry of $(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{S}_{t}\in\mathbb{R}^{m\times n}$ being
$\left[(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{S}_{t}\right]_{j,j}=\frac{[\tilde{S}_{t}]_{j,j}}{\sqrt{[\tilde{S}_{t}]_{j,j}^{2}+\epsilon}},\quad\text{for all $j\leq\min\left\{m,n\right\}$}$
We can see that if $[\tilde{S}_{t}]_{j,j}=0$ then the above gives 00. Thus when taking $\epsilon\to 0$, ASGO is
$W_{t+1}=W_{t}-\eta_{t}U_{t}V_{t}^{\top},$
where $U_{t}=\tilde{U}_{t}[:,:r_{t}]$ and $V_{t}=\tilde{V}_{t}[:,:r_{t}]$ with $r_{t}$ as the rank of $G_{t}$. This is the same as Muon. ∎
We then prove a nonconvex convergence result for Muon in the deterministic case with $\mu=0$.
Theorem 4 (Nonconvex convergence of Muon).
We consider Muon presented in (10) with $\mu=0$. In the deterministic case, i.e. no gradient noise, if we assume that there exists a lower bound $f^{*}$ such that $f(W)\geq f^{*}$ for all $W$ and Assumption 2, by taking $\eta_{t}\equiv\eta=\sqrt{\frac{2(f(W_{0})-f^{*})}{\left\|L\right\|_{*}T}}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f(W_{t})\right\|_{*}\leq\sqrt{\frac{\left\|L\right\|_{*}(f(W_{0})-f^{*}))}{2T}}.$
Proof.
Since we assume the deterministic setting, we have $G_{t}=\nabla f(W_{t})$. Also, since we assume $\mu=0$, we have $U_{t},V_{t}$ are the left and right singular vectors of $G_{t}$ such that $G_{t}=U_{t}S_{t}V_{t}^{\top}$ Then based on the smoothness assumption, it holds that
$f(W_{t+1})\leq$ $f(W_{t})+\left\langle\nabla f(W_{t}),W_{t+1}-W_{t}\right\rangle+\frac{1}{2}{\rm tr}\left((W_{t+1}-W_{t})^{\top}L(W_{t+1}-W_{t})\right)$
$=$ $f(W_{t})-\eta_{t}{\rm tr}\left(G_{t}^{\top}U_{t}V_{t}^{\top}\right)+\frac{\eta_{t}^{2}}{2}{\rm tr}\left((U_{t}V_{t}^{\top})^{\top}L(U_{t}V_{t}^{\top})\right)$
$=$ $f(W_{t})-\eta_{t}{\rm tr}\left(V_{t}S_{t}U_{t}^{\top}U_{t}V_{t}^{\top}\right)+\frac{\eta_{t}^{2}}{2}{\rm tr}\left(V_{t}U_{t}^{\top}LU_{t}V_{t}^{\top}\right)$
$=$ $f(W_{t})-\eta_{t}{\rm tr}\left(V_{t}S_{t}V_{t}^{\top}\right)+\frac{\eta_{t}^{2}}{2}{\rm tr}\left(U_{t}^{\top}LU_{t}V_{t}^{\top}V_{t}\right)$
$=$ $f(W_{t})-\eta_{t}{\rm tr}\left(S_{t}V_{t}^{\top}V_{t}\right)+\frac{\eta_{t}^{2}}{2}{\rm tr}\left(LU_{t}U_{t}^{\top}\right)$
$\leq$ $f(W_{t})-\eta_{t}\left\|G_{t}\right\|_{*}+\frac{\eta_{t}^{2}}{2}\left\|L\right\|_{*},$
where the last inequality is because of the fact that $UU^{\top}\preceq I$ and Lemma 1. Then by summing up over $t$ and rearrangement, we can obtain that
$\sum_{t=0}^{T-1}\eta_{t}\left\|G_{t}\right\|\leq f(W_{0})-f(W_{T})+\sum_{t=0}^{T-1}\frac{\eta_{t}^{2}}{2}\left\|L\right\|_{*}.$
Then we take $\eta_{t}\equiv\eta$ to obtain that
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f(W_{t})\right\|\leq\frac{f(W_{0})-f^{*}}{\eta T}+\frac{\eta}{2}\left\|L\right\|_{*}.$
Setting $\eta=\sqrt{\frac{2(f(W_{0})-f^{*})}{\left\|L\right\|_{*}T}}$ finishes the proof. ∎